OpenStudy (anonymous):

Simplify the expression. (-2 - c)(-1)

OpenStudy (anonymous):

@ganeshie8

OpenStudy (sleepyhead314):

distribute the (-1) so (-2)(-1) + (-c)(-1) = ?

OpenStudy (party_girl):

-(-c-2)

OpenStudy (anonymous):

I don't know how to get the answer to (-2)(-1) + (-c)(-1) =

OpenStudy (anonymous):

Can you possibly explain it, step by step?

OpenStudy (sleepyhead314):

distributive property a(b-c) = a*b + a*(-c)

OpenStudy (anonymous):



OpenStudy (anonymous):

then minus * minus= positve hence( -2)(-1)=2